The perimeter of a rectangular garden is 26 feet. The length is 7 feet longer than the width. What is the length of the garden?
Answer by Cromlix(4381)

Hi there,
Make width = x
Length = 7 + x
Perimeter = 2 x Width + 2 x Length
26 = 2(x) + 2(7 + x)
26 = 2x + 14 + 2x
Collect like terms
2x + 2x = 26 - 14
4x = 12
x = 3
Width = 3 feet
Length = 10 feet.
Hope this helps :-)